OpenStudy (anonymous):

Which number equals (2)^-4

OpenStudy (anonymous):

A number to the power of a negative number is equal to 1 over that number. In general terms, a^(-1) = 1/a. Now if that index number is another number (e.g 2) then that number goes at the bottom of the fraction. In general terms, a^(-b) = 1 / (a^b) In your case, you have 2^(-4). So that is 1/ (2^4). 2^4 = 16 so your answer is 1/16

OpenStudy (anonymous):

\[\large (a)^{-x} = \frac{1}{a^x}\]

OpenStudy (anonymous):

@benji13 Got or not?

OpenStudy (anonymous):

yes i goy @waterineyes

OpenStudy (anonymous):

Good then..

OpenStudy (anonymous):

yea